OpenStudy (anonymous):

let f(x) =3x^3 -5x^2+4x-6 and g(x) = -8x^2-7x +12,evaluate (f.g)(1/2) a.13/8 b.13/6 c.-13/4 d.-13/8

OpenStudy (anonymous):

lag dont lag this is 2nd to the last question

OpenStudy (anonymous):

im getting error now when it comes to 4x-6 short term memory

OpenStudy (anonymous):

13/6 letter b

OpenStudy (anonymous):

ops w8 wrong theres 1/2

OpenStudy (anonymous):

13/6 sir?

OpenStudy (anonymous):

\[27x-10x+4x =13 16x-7=9 then 9-12=-3 -6-3 =3 f.g(13/3) (1/2)=13/6\]

OpenStudy (anonymous):

ops wrong 1316x =only 13x then 16x-7

OpenStudy (anonymous):

is it b or d

OpenStudy (anonymous):

what is your answer sir?

OpenStudy (ash2326):

we have to find f.g(1/2) let's find f(x).g(x) \[f(x).g(x)=(3x^3-5x^2+4x-6)(-8x^2-7x+12)\] substitute x=1/2 \[f.g(1/2)=(3/8-5/4+4/2-6)(-8/4-7/2+12)\] on solving we get f.g(1/2)=-507/16 not matching any option sorry i'd to go somehere in the middle

OpenStudy (ash2326):

*somewhere

OpenStudy (anonymous):

it will be (-8x-7x)^2 = 8

OpenStudy (anonymous):

-8 or 8

OpenStudy (ash2326):

how 'd you get this?

OpenStudy (anonymous):

im checking my notebook example of my prof XD its not finalized XD

OpenStudy (anonymous):

sqrt of 507 = 13sqrt3

OpenStudy (anonymous):

letter D final not sure